Stella’s Catering is planning a wedding
reception. The bride and groom would like to serve
a nut mixture containing 25% peanuts. Stella has
available mixtures that are either 40% or 10% peanuts.
How much of each type should be mixed to get a 10-lb
mixture that is 25% peanuts?
: Stella’s Catering is planning a wedding
reception. The bride and groom would like to serve
a nut mixture containing 25% peanuts. Stella has
available mixtures that are either 40% or 10% peanuts.
How much of each type should be mixed to get a 10-lb
mixture that is 25% peanuts?

Answer by stanbon(18991)

The bride and groom would like to serve
a nut mixture containing 25% peanuts.
Stella has available mixtures that are
either 40% or 10% peanuts.
How much of each type should be mixed to get a 10-lb
mixture that is 25% peanuts?
-------------------------------------------------
Amt. of 40% mix = "x" lbs. ; Amt of peanuts in it = 0.40x lbs.
--------------------------
Amt. of 10% mix = "10-x" lbs; Amt. of peanuts in it = 0.10(10-x) = 1-0.1x lbs
----------------------------
Amt of 25% mis = 10 lbs ; Amt. of peanuts in it = 0.25(10) = 2.5 lbs
===========================
EQUATION:
peanuts + peanuts = peanuts
0.40x + 1 - 0.1x = 2.5
0.3x = 1.5
x = 5 lbs (amt. of 40% nuts in the final mixture)
10-x = 5 lbs (amt of 10% nuts in the final mixture)
=============================
Cheers,
Stan H.